# U-DECN: End-to-End Underwater Object Detection ConvNet with Improved DeNoising Training

### v1

@@ 3.3 Separate Contrastive DeNoising Forward
The Contrastive DeNoising (CDN) training of DINO [28] is very effective in stabilizing training and accelerating convergence. It constructs multiple sets of positive and negative queries for object class and boxes, effectively inhibiting duplicate predictions and selection of anchor predictions (in mixed query selection) further from the ground truth (GT) boxes. The CDN queries are concatenated with object queries and then interact with encoder feature embeddings by cross-attention. However, the decoder of U-DECN is ConvNet and cannot interact with encoder features in this way. To address this issue, we propose a separate contrastive denoising forward method to perform the denoising training of the ConvNet architecture in Fig. 2. During training, the ConvNet decoder interacts with encoder features twice, separately for object queries and CDN queries. Our separate contrastive denoising forward avoids using the attention mask, as the separate forward prevents object queries and CDN queries from seeing each other, preventing information leakage in the SIM. In addition, in our method, the number of CDN queries must be the same as the object queries to achieve consistency in information interaction based on the ConvNet decoder. Meanwhile, the positional queries in CDN constructed by GT boxes are also uncertainly arranged, leading to the query feature misalignment issue to CDN queries as the same, details in Sec. 3.4.
3.4 Deformable Convolution in SIM
As mentioned above, the positional queries constructed using selected top-K features and GT boxes both have the query feature misalignment issue. In CIM, due to the strong positional inductive bias of ConvNets, queries $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$ could only interact with the encoder feature corresponding to their tensor positions. The misalignment between uncertainly arranged positional queries and encoder features results in low interaction efficiency between them, which in turn reduces training convergence speed and model performance, analysis from Tab. 4 and Fig. 5. In SIM, even if it uses a $9\times 9$ large kernel depth-wise convolution for self-interaction between queries, which has a large receptive field, it is difficult to correct the position of uncertainly arranged queries. To address this, we replace the $9\times 9$ large kernel depth-wise convolution in SIM with a $9\times 9$ large kernel deformable convolution (DCNv3 [25]), as shown in Fig. 3. The deformable convolution achieves deformable operations by predicting the offset of each feature point in the feature map, and the predicted feature point offset is equivalent to using self-attention to correct the feature position. Based on the above analysis, deformable convolution can also correct the position of uncertainly arranged positional queries in tensor $\bm{q}$ by learning the bounding box position information in priori positional queries to improve the efficiency of interaction with encoder features.
3.5 Hybrid Encoder
The ConvNet encoder-decoder architecture of DECO [4] can only process single-level features, and cannot perform the interaction of multi-level features like the detector transformer [31], resulting in a lack of multi-scale feature information and reducing the performance of the detector for objects on different scales. The simultaneous intra-scale and cross-scale feature interaction in the transformer encoder has been proven to be inefficient cf. RT-DETR [30], resulting in a computational bottleneck. RT-DETR proposes an efficient hybrid encoder, consisting of the Attention-based Intra-scale Feature Interaction (AIFI) and the CNN-based Cross-scale Feature Fusion (CCFF) modules. Specifically, AIFI further reduces the computational cost by performing the intra-scale interaction only on the highest-level feature. Then, CCFF fuses the highest-level feature and other low-level features and outputs the fusion features of the corresponding levels. To introduce multi-scale feature information, we also use the hybrid encoder architecture, as shown in Fig. 2, where AIFI i.e. our ConvNet encoder. We take the highest-level feature of CCFF for decoder, since our ConvNet decoder could still only process single-level features.
3.6 Underwater Color DeNoising Query
@@
DINO (1scale) [28] 12 41 92 27 58.5 80.6 66.5 46.6 60.6 56.7
U-DECN 50 82 170 29 61.4(+2.9) 84.7 70.6 52.1 63.8 59.0
U-DECN 72 82 170 29 63.3(+4.8) 85.0 72.6 51.0 65.3 61.8
U-DECN 100 82 170 29 64.0(+5.5) 84.1 72.8 48.2 65.4 63.0
Table 1: Comparisons of our U-DECN with other end-to-end detectors. The FPS is reported on an NVIDIA A6000 GPU.
@@
DINO (4scale) [28] 47 262 4 63.5
Deformable DETR (4scale) [31] 41 196 4 64.0
U-DECN (1scale) 82 170 21(+17) 64.0
Table 2: Comparisons of our U-DECN with 4-scale detectors. The $\textbf{FPS}_{trt}$ is measured on the NVIDIA AGX Orin by TensorRT FP16.
We rethink the effect of underwater color cast noise and consider the color cast noise as color fog. However, there is a difference between this color fog and normal white fog. White fog on land blocks the light reflected from objects, resulting in a lack of object feature information. Color fog is caused by the absorption of light by water media. Although the light reflected from objects is also absorbed causing color cast, the color cast results in a bias rather than a lack of object feature information. Therefore, in terms of noise generalization, we propose an underwater Color DeNoising (ColorDN) query to improve the generalization of the model for the biased object feature information by different color cast noise, as shown in Figs. 1 and 2. We simply define the RGB values of color fog in underwater images as the mean of the three RGB channels in the image, and encode it as a ColorDN query using a liner layer, which is as follows:
$\bm{q}_{colordn}=\mathrm{Liner}\left(\mu_{c}(\bm{x}_{input})\right)$ (1)
$\bm{q}_{content}=\bm{q}_{content}+\bm{q}_{colordn}$ (2)
where for $N$ different content queries (the same for CDN queries), we use the tensor repeat operation to extend the ColorDN query to make their quantity consistent.
We introduce priori color cast noise information into content queries to enable the model to query the corresponding noise information, thereby generalizing the bias of object feature information. The underwater ColorDN query improves the generalization of the detector to color cast noise in a simple and implicit way.
@@ 4 Experiments
Dataset and evaluation metrics. We conduct experiments on the underwater dataset DUO [14], which includes the holothurian, echinus, scallop, and starfish classes. The DUO dataset consists of the training set (6671 images) and the test set ($1$ images) for training and validation. We follow the standard COCO [12] evaluation protocol and report the Average Precision (AP).
@@
Implementation details. Our implementations are based on MMDetection toolbox [2]. All object detection models are trained on 2 NVIDIA A6000 GPUs with batch size 4 per GPU. The training hyper-parameters almost follow DINO [28], we train the proposed U-DECN using AdamW [19] optimizer, with weight decay of $1e^{-4}$ and initial learning rates (lr) as $2e^{-4}$ and $2e^{-5}$ for the encoder-decoder and backbone, respectively. We drop lr at the 40-th, 60-th, and 80-th epoch by multiplying 0.1 for the 50, 72, and 100 epoch settings. We follow the same augmentation scheme as DINO, which includes randomly resizing the input image such that its short side is between 480 and 800 pixels and its long size is at most 1333 pixels, randomly flipping the training image with probability 0.5 on horizontal direction, randomly cropping the training image with probability 0.5 to a random rectangular patch.
Model DUO DUO(H=42.5) DUO(H=67) DUO(H=91.5)
DECO [4] 46.0 27.1 38.1 29.8
DETR [1] 48.4 31.9 39.9 31.3
Deformable DETR (1scale) [31] 52.6 32.9 43.4 32.0
DAB-DETR [15] 56.5 36.1 46.5 37.2
Conditional-DETR [20] 56.8 37.3 47.3 36.8
DINO (1scale) [28] 58.5 37.5 48.8 38.0
U-DECN w/o ColorDN 61.8 35.4 51.0 37.6
U-DECN 63.3 38.5 52.3 38.2
4.1 Comparisons with State-of-the-arts
@@
Comparisons with End-to-End Detectors. We evaluate the proposed U-DECN on DUO [14] benchmark and compare it with recent competitive object detectors, including DECO [4], DETR [1], Deformable DETR [31], DAB-DETR [15], Conditional-DETR [20], and DINO [28], whose backbones are ResNet-50 [9]. Experimental results in terms of detection AP and GFLOPs/FPS are shown in Tabs. 1 and 2, where the GFLOPs and FPS are measured on the input size of (1333, 800). In 1-scale comparisons Tab. 1, our U-DECN achieves a significant performance improvement of 15.4-18.0 AP compared to the DECO based on the same convolutional encoder-decoder. The training convergence curves also validate the effectiveness of our methods in improving both convergence speed and performance, as in Fig. 4. Meanwhile, our U-DECN obtains 61.4, 63.3, 64.0 AP at 50, 72, 100 epochs and are 2.9, 4.8, 5.5 AP better than DINO. Although U-DECN has a higher number of parameters and GFLOPs, it still has a similar running speed FPS. In 4-scale comparisons Tab. 2, our 1-scale U-DECN achieves 64.0 AP and has similar performance compared to 4-scale Deformable DETR (64.0 AP) and DINO (63.5 AP). We deploy detection models on the NVIDIA AGX Orin by TensorRT FP16. Compared to DINO and Deformable DETR, U-DECN significantly improves the speed by 5 times (21 FPS vs 4 FPS), which meets the real-time requirements for deployment on underwater vehicle platforms.
@@
Comparisons on Different Underwater Color Casts. We evaluate the performance of models on DUO [14] with different underwater color casts to validate the generalization performance in Tab. 3. We select three hue values at equidistant intervals in the underwater color cast range [18] for the test (models are all trained on raw DUO). Our U-DECN outperforms others for 38.5, 52.3, 38.2 AP on 42.5, 67, 91.5 H, which is a significant improvement in generalization performance compared to it w/o ColorDN.
4.2 Ablation Study
Ablation Study on U-DECN. Tab. 4 shows the improvement of TS, CDN, DCN, Hybrid, and ColorDN methods, which significantly improve the performance of the U-DECN detector. Of these, the TS method achieves the highest performance improvement of 8.0 AP. This ablation study demonstrates the effectiveness of the designed methods on ConvNet encoder-decoder architecture. While, the DCN method significantly improves both performance and convergence speed in Fig. 5, which effectively improves the query feature misalignment issue as discussed in Sec. 3.4.
@@ 4.2 Ablation Study
Ablation Study on DCN in SIM. We compare different offset scale and groups settings of DCNv3 [25] in SIM, as shown in Tab. 5. When the offset scale is set to 1.0, U-DECN achieves a better performance of 60.7 AP, but 480 groups make the model a huge number of parameters 397 M. Thus, we ablate different groups and find that U-DECN achieves the best performance of 60.8 AP at 30 FPS with groups set to 16, and the number of model parameters drops significantly to 72 M.
@@
100 (10$\times$10) 63.3
100 (10$\times$10) 60.3
225 (15$\times$15) 62.0
225 (15$\times$15) 59.5
400 (20$\times$20) 58.4 (†58.4)
625 (25$\times$25) 58.6 (†58.7)
900 (30$\times$30) 58.2 (†58.3)
[…]
Ablation Study on Different Numbers of Queries. We ablate different numbers of queries using 100 queries in the test, where the query shape is that $h_{q}:w_{q}=1:1$, as shown in Tab. 6. Meanwhile, the multi-level features from the hybrid encoder can be selected to initialize the positional queries. For over 400 queries, we choose only mlvl feats because the number of single-level feature patches is not sufficient to initialize positional queries. Our U-DECN achieves the best performance of 63.3 AP using 100 queries w/o mlvl feats. The performance drops significantly w/ mlvl feats, which is caused by the decoder interacting with single-level value using multi-level query. Despite the multi-scale feature fusion, the problem of cross-level feature interaction in query and value still exists. As the number of queries gradually increases, the performance of the model gradually decreases (a tiny increase at 300 queries in the test), because the large number of queries makes learning difficult for the ConvNet encoder-decoder architecture (while the transformer encoder-decoder architecture does not suffer from this problem).
4.3 Visual Analysis
We visualize the detection results in Fig. 6. The visual detection results demonstrate that our U-DECN is more robust than other detectors in underwater environments with severe color cast. It can be observed that U-DECN performs better on objects in underwater situations, such as dense, occluded, and noisy scenarios. These results highlight the robustness and adaptability of U-DECN to the unique challenges of underwater detection.
@@ 5 Conclusion
In this paper, we propose an end-to-end underwater object detector called U-DECN based on DECO, with ConvNet encoder-decoder architecture. We integrate advanced technologies from DETR variants into DECO and design optimization methods specifically for the ConvNet architecture, including Separate Contrastive DeNoising Forward and Deformable Convolution in SIM. Furthermore, we propose an underwater color denoising query that improves the generalization for underwater color cast noise. Extensive experiments and ablation studies validate the effectiveness of our methods in improving convergence speed, performance, running speed, and deployment efficiency (especially on NVIDIA AGX Orin by TensorRT FP16). And U-DECN outperforms the other state-of-the-art end-to-end object detectors on DUO.
However, U-DECN still has some limitations. U-DECN is still unable to use the multi-level feature as the value of the decoder due to its ConvNet encoder-decoder architecture, which limits its performance on multi-scale object detection. The increase in the number of queries makes model learning difficult and reduces the performance of the detector. Moreover, our U-DECN can only generalize underwater color cast noise, but other underwater noise also has a detrimental effect on the detector. The above problems still remain for future exploration. We hope that these problems will be addressed in future work.
Acknowledgments. This research is funded by the National Natural Science Foundation of China, grant number 52371350, by the National Key Laboratory Foundation of Autonomous Marine Vehicle Technology, grant number 2024-HYHXQ-WDZC03, by the Natural Science Foundation of Hainan Province, grant number 2021JJLH0002.

### v2

@@ 3.4 Separate Contrastive DeNoising Forward
Table 1: Comparisons of our U-DECN with other end-to-end detectors on DUO. The FPS is reported on an NVIDIA A6000 GPU.
@@
DINO [30] 12 47 262 15 63.5 82.1 69.9 48.1 65.4 61.9
Deformable DETR [34] 50 41 196 18 64.0 85.5 72.1 50.2 66.1 62.2
DECO [5] 150 56 105 41 46.0 72.4 51.2 33.8 46.5 46.6
U-DECN 50 82 170 29 61.4(+15.4) 84.7 70.6 52.1 63.8 59.0
U-DECN 72 82 170 29 63.3(+17.3) 85.0 72.6 51.0 65.3 61.8
U-DECN 100 82 170 29 64.0(+18.0) 84.1 72.8 48.2 65.4 63.0
Table 2: Comparisons of our U-DECN in terms of deployment. The $\textbf{FPS}_{trt}$ is measured on the NVIDIA AGX Orin by TensorRT FP16.
@@
DINO [30] 47 262 4 63.5
Deformable DETR [34] 41 196 4 64.0
U-DECN 82 170 21(+17) 64.0
The Contrastive DeNoising (CDN) training of DINO [30] is very effective in stabilizing training and accelerating convergence. It constructs multiple sets of positive and negative queries for object class and boxes, effectively inhibiting duplicate predictions and selection of anchor predictions (in mixed query selection) further from the ground truth (GT) boxes. The CDN queries are concatenated with object queries and then interact with encoder feature embeddings by cross-attention. However, the decoder of U-DECN is ConvNet and cannot interact with encoder feature embeddings in this way, it requires the queries with the fixed tensor shape and quantity. To address this issue, we propose a separate contrastive denoising forward method to perform the denoising training of the ConvNet architecture as shown in Fig. 2. During training, the ConvNet decoder interacts with encoder feature embeddings $\bm{x}_{f}$ (which is from the hybrid encoder) twice, separately for object queries $\bm{q}_{obj}$ and CDN queries $\bm{q}_{cdn}$. The quantity of CDN queries must be the same as the object queries, and they must meet the settings of the ConvNet decoder for the query quantity $N$ and its tensor shape $d\times h_{q}\times w_{q}$. We use the method in DINO to construct CDN queries more than the set quantity, taking the first $N$ to reshape the queries to $d\times h_{q}\times w_{q}$ tensor shape. The loss $\mathcal{L}$ of our U-DECN is expressed as follows:
$\mathcal{L}_{hungarian}$ $=\mathrm{Loss}\left(\mathrm{Decoder}\left(\bm{x}_{f},\bm{q}_{obj}\right),GT\right)$ (3)
$\mathcal{L}_{cdn}$ $=\mathrm{Loss}_{cdn}\left(\mathrm{Decoder}\left(\bm{x}_{f},\bm{q}_{cdn}\right),GT_{cdn}\right)$ (4)
$\mathcal{L}$ $=\mathcal{L}_{hungarian}+\mathcal{L}_{cdn}$ (5)
where $\mathcal{L}_{hungarian}$ and $\mathcal{L}_{cdn}$ represent the object detection loss with CDN loss, which is the same as those in DINO. In our separate contrastive denoising forward method, object queries $\bm{q}_{obj}$ and CDN queries $\bm{q}_{cdn}$ do not interact, which also prevents information leakage between object queries and CDN queries. Meanwhile, the positional queries in CDN constructed by GT boxes are also uncertainly arranged, leading to the query feature misalignment issue to CDN queries as the same. The issue can also be addressed by the Deformable Convolution in SIM method.
@@ 4.1 Dataset and Evaluation Metrics
We conduct experiments on the underwater dataset DUO [17] and RUOD [8]. All models in the experiments are trained on either DUO or RUOD and tested on the corresponding dataset. The DUO dataset consists of the training set (6671 images) and the test set (1111 images) for training and validation. It includes 4 categories of underwater organisms. The RUOD dataset consists of the training set (9800 images) and the test set (4200 images) for training and validation. It contains various underwater scenarios and consists of 10 categories. We follow the standard COCO [15] evaluation protocol and report the Average Precision (AP).
4.2 Implementation Details
@@
Our implementations are based on MMDetection toolbox [3]. All object detection models are trained on 2 NVIDIA A6000 GPUs with batch size 4 per GPU. The training hyper-parameters almost follow DINO [30], we train the proposed U-DECN using AdamW [21] optimizer, with weight decay of $1e^{-4}$ and initial learning rates (lr) as $2e^{-4}$ and $2e^{-5}$ for the encoder-decoder and backbone, respectively. We drop lr at the 40-th, 60-th, and 80-th epoch by multiplying 0.1 for the 50, 72, and 100 epoch settings. We follow the same augmentation scheme as DINO, which includes randomly resizing the input image such that its short side is between 480 and 800 pixels and its long size is at most 1333 pixels, randomly flipping the training image with probability 0.5 on horizontal direction, randomly cropping the training image with probability 0.5 to a random rectangular patch.
Table 3: Comparisons of our U-DECN with other end-to-end detectors on RUOD.
Model #Params (M) GFLOPs AP $\textbf{AP}_{50}$ $\textbf{AP}_{75}$
DETR [1] 42 97 47.0 77.7 48.6
DAB-DETR [18] 44 98 52.1 83.0 56.5
Conditional-DETR [22] 43 97 52.2 82.3 56.7
DINO [30] 47 262 57.8 82.3 62.9
Deformable DETR [34] 41 196 57.8 85.5 62.9
DECO [5] 56 105 53.1 82.5 57.1
U-DECN 82 170 55.8(+2.7) 85.1 61.6
U-DECN 82 170 57.5(+4.4) 85.4 63.4
U-DECN 82 170 58.1(+5.0) 84.5 63.6
4.3 Quantitative and Qualitative Comparisons
4.3.1 Comparisons with End-to-End Detectors
@@
We evaluate the proposed U-DECN on DUO [17] and RUOD [8] benchmark and compare it with recent competitive object detectors, including DECO [5], DETR [1], Deformable DETR [34], DAB-DETR [18], Conditional-DETR [22], and DINO [30], whose backbones are ResNet-50 [11]. Experimental results in terms of detection APs and GFLOPs/FPS are shown in Tables 1, 3 and 2, where the GFLOPs and FPS are measured on the input size of (1333, 800). In Tables 1 and 3, our U-DECN achieves a significant performance improvement compared to the DECO based on the same convolutional encoder-decoder architecture. Compared to DECO, which is also a ConvNet architecture, the training convergence curves also validate the effectiveness of our methods in improving both the convergence speed and the performance of DECO, as shown in Fig. 4. In Table 2, our U-DECN achieves 64.0 AP and has similar performance compared to Deformable DETR (64.0 AP) and DINO (63.5 AP). We deploy detection models on the NVIDIA AGX Orin by TensorRT FP16. Compared to DINO and Deformable DETR, U-DECN significantly improves the speed by 5 times (21 FPS vs 4 FPS), which meets the real-time requirements for deployment on underwater vehicle platforms. Under similar performance conditions, our U-DECN has a smaller GFLOPs and larger FPS than Deformable DETR and DINO, despite having a larger number of parameters. In conclusion, our method exhibits exemplary running efficiency and convergence speed, underscoring its profound potential for real-world underwater applications.
4.3.2 Comparisons on Different Underwater Color Casts
We evaluate the performance of models on DUO [17] with different underwater color casts to validate the generalization performance in Table 5. We select three hue values at equidistant intervals in the underwater color cast range [20] for the test (models are all trained on raw DUO). The different H represents the color cast in various color ranges underwater. Our U-DECN outperforms others for 39.4, 53.2, 39.1 AP on 42.5, 67, 91.5 H, which is a significant improvement in generalization performance compared to it w/o ColorDN.
@@ 4.3.2 Comparisons on Different Underwater Color Casts
4.3.3 Visual Comparisons
We visualize the detection results on images with different color casts in Fig. 5. It can be observed that U-DECN performs better on objects in underwater situations compared to other detectors, such as dense, occluded, and noisy scenarios. The visual detection results demonstrate that our U-DECN is more robust than other competitive detectors in underwater environments with severe color cast. These results highlight the robustness and adaptability of U-DECN to the unique challenges of underwater detection.
@@ 4.3.3 Visual Comparisons
Model DUO DUO(H=42.5) DUO(H=67) DUO(H=91.5)
DECO [5] 46.0 27.1 38.1 29.8
DETR [1] 48.4 31.9 39.9 31.3
DAB-DETR [18] 56.5 36.1 46.5 37.2
Conditional-DETR [22] 56.8 37.3 47.3 36.8
DINO [30] 63.5 38.2 52.9 38.6
Deformable DETR [34] 64.0 38.8 52.7 38.0
U-DECN w/o ColorDN 62.5 36.2 51.9 38.5
U-DECN 64.0 39.4 53.2 39.1
4.4 Ablation Studies
4.4.1 Ablation Study on U-DECN
Table 4 shows the improvement of TS, CDN, DCN, Hybrid, and ColorDN methods, which significantly improve the performance of the U-DECN detector. Of these, the TS method achieves the highest performance improvement of 8.0 AP. This ablation study demonstrates the effectiveness of the designed methods on ConvNet encoder-decoder architecture. While, the DCN method significantly improves both performance and convergence speed in Fig. 6, which effectively improves the query feature misalignment issue as discussed in Section 3.2.
@@ 4.4.1 Ablation Study on U-DECN
Table 7: Results of the ablation study for different numbers of queries. The mlvl feats represents using multi-level features to initialize the positional queries. $\dagger$} represents using 300 queries in the test, while others are 100 queries.
@@
100 ($\times10$) 63.3
100 ($\times10$) 60.3
225 ($\times15$) 62.0
225 ($\times15$) 59.5
400 ($\times20$) 58.4 (†58.4)
625 ($\times25$) 58.6 (†58.7)
900 ($\times30$) 58.2 (†58.3)
4.4.2 Ablation Study on DCN in SIM
We compare different offset scale and groups settings of DCNv3 [28] in SIM, as shown in Table 6. When the offset scale is set to 1.0, U-DECN achieves a better performance of 60.7 AP, but 480 groups make the model a huge number of parameters 397 M. Thus, we ablate different groups and find that U-DECN achieves the best performance of 60.8 AP at 30 FPS with groups set to 16, and the number of model parameters drops significantly to 72 M.
4.4.3 Ablation Study on Different Numbers of Queries
We ablate different numbers of queries using 100 queries in the test, where the query shape is that $h_{q}:w_{q}=1:1$, as shown in Table 7. Meanwhile, the multi-level features from the hybrid encoder can be selected to initialize the positional queries. For over 400 queries, we choose only mlvl feats because the number of single-level feature patches is not sufficient to initialize positional queries. Our U-DECN achieves the best performance of 63.3 AP using 100 queries w/o mlvl feats. The performance drops significantly w/ mlvl feats, which is caused by the decoder interacting with single-level value using multi-level query. Despite the multi-scale feature fusion, the problem of cross-level feature interaction in query and value still exists. As the number of queries gradually increases, the performance of the model gradually decreases (a tiny increase at 300 queries in the test), because the large number of queries makes learning difficult for the ConvNet encoder-decoder architecture (while the transformer encoder-decoder architecture does not suffer from this problem).
@@ 5 Conclusion
In this paper, we propose an end-to-end underwater object detector called U-DECN based on DECO, with ConvNet encoder-decoder architecture. We integrate advanced technologies from DETR variants into DECO and design optimization methods specifically for the ConvNet architecture, including Deformable Convolution in SIM and Separate Contrastive DeNoising Forward. Furthermore, we propose an Underwater Color DeNoising Query method that improves the generalization for underwater color cast noise. Extensive experiments and ablation studies validate the effectiveness of our methods in improving convergence speed, performance, running speed, and deployment efficiency (especially on NVIDIA AGX Orin by TensorRT FP16). And U-DECN outperforms the other state-of-the-art end-to-end object detectors on DUO and RUOD. U-DECN is particularly suitable for real-time deployment on embedded platforms with limited processing power, such as unmanned underwater vehicles (UUVs).
However, U-DECN still has some limitations. The increase in the number of queries makes model learning difficult and reduces the performance of the detector. Moreover, our U-DECN can only generalize underwater color cast noise, but other underwater noise also has a detrimental effect on the detector, such as suspended solids in water, uneven underwater lighting conditions, refraction and reflection of underwater light, and artificial underwater light sources. The above problems still remain for future exploration. We hope that these problems will be addressed in future work.
Acknowledgments. This research is funded by the National Natural Science Foundation of China, grant number 52371350, by the National Key Research and Development Program of China, grant number 2023YFC2809104, and by the National Key Laboratory Foundation of Autonomous Marine Vehicle Technology, grant number 2024-HYHXQ-WDZC03.